# Data Collection with Non-Uniform Axial Power for Phase II of the OECD/NEA AI/ML Critical Heat Flux Benchmark

### v1

@@
Data Collection with Non-Uniform Axial Power for Phase II of the OECD/NEA AI/ML Critical Heat Flux Benchmark
@@ 1 Introduction
Critical heat flux (CHF) is a key thermal-hydraulic phenomenon in which the two-phase coolant’s heat-transfer capability drops abruptly, establishing the thermal-design limit, defined as the departure from nucleate boiling (DNB) in pressurized water reactors (PWRs) and as the critical-power condition in boiling water reactors (BWRs) [1]. Therefore, accurate prediction of CHF is essential for setting safe operating margins and maximizing reactor power output without risking fuel damage. However, CHF modeling remains challenging due to the complex coupling of local two-phase hydrodynamics, surface characteristics, transient thermal boundary-layer development, and geometric influences, factors that vary widely with pressure, flow rate, and heating profile [2, 3].
Traditional CHF correlations often rely on simplified mechanistic assumptions, parametric relationships or look-up tables (LUT), which can limit their accuracy and adaptability under novel or strongly non-uniform heating conditions [4, 5]. Machine learning (ML) methods, by contrast, can learn high-dimensional patterns directly from large datasets and have demonstrated improved accuracy and robustness in CHF prediction [6]. However, many ML-based CHF studies have been carried out independently as proof-of-concept demonstrations on limited or proprietary datasets, underscoring the need for coordinated benchmarking efforts. To address this, the OECD/NEA Task Force on Artificial Intelligence and Machine Learning for Scientific Computing in Nuclear Engineering organized a CHF benchmark to provide common datasets, blind-test partitions, and reproducible evaluation protocols for evaluating ML-based CHF predictors [7].
Phase I of the CHF benchmark, as specified in [7], comprised four tasks: (1) dimensionality analysis for feature selection and extraction; (2) development and optimization of ML-based CHF regression models based on the Nuclear Regulatory Commision (NRC) database; (3) rigorous model evaluation and uncertainty quantification (UQ); and (4) independent blind testing on withheld data. While Phase I demonstrated the feasibility of AI/ML approaches for CHF prediction under uniform heating, it remained limited by the absence of spatially varying power profiles. Phase II will extend the CHF benchmark to include more complex geometries and boundary conditions, to be used along with transfer learning techniques advanced UQ, and design optimization, all of which require a richer, more diverse CHF dataset.
To support these Phase II tasks, Kansas State University, in collaboration with Westinghouse Electric Sweden, collected additional CHF measurements in tubes with uniform and non-uniform axial power distributions within the ME 535 “Interdisciplinary Industrial Design Project II” course. This paper reports on those activities and describes the newly digitized uniform and non-uniform CHF datasets; more details can be found in [8].
@@ 2.1 Data Source and Characteristics
The primary source of the experimental data collected comes from the KAERI technical report [9]. The report investigates how axial heat flux distribution affects the occurrence of CHF in boiling systems. The study includes an extensive set of experimental CHF measurements under both uniform and non-uniform heating conditions, providing valuable data for understanding and modeling CHF behavior in nuclear thermal-hydraulic systems.
Overall, $1539$ CHF test cases were collected/digitized from [9], comprising $651$ uniform-heating measurements and $888$ non-uniform heating cases, summarized in Table 1. The experimental setups used water coolant in round tubes with internal diameter of 5.445.445.445.4428.328.328.328.3 mm and heated lengths of 0.0610.0610.0610.0617.07.07.07.0 m.
@@
Uniform 0.43 – 18 335.0 – 9561.9 $-1.461$ – 0.890
Non-uniform 0.43 – 18 328.2 – 8916.0 $-1.352$ – 0.804
@@ 2.2 Data Collection Workflow
    Energy balance check: The integrated axial flux was compared to published total power; cases with $>$2% discrepancy were reprocessed.
@@
For the uniform cases, inlet subcooling and outlet quality are reported in [9]. The remaining inlet and outlet parameters were calculated using thermodynamic properties from [12]. Unlike the non-uniform cases, the uniform data did not use a detailed 40-node mesh; instead, only inlet and outlet values of local power and temperature/equilibrium quality were recorded. In the non-uniform dataset, the inlet temperature is provided and, combined with the specified heat flux distribution, was used to compute all thermodynamic parameters along the heated tube on a uniform 40-node mesh.
@@ 2.3 Dataset Structure
The quantities were saved in separate (uniform and non-uniform) XML files which are easily readable by computer scripts, see Table 2 that describes the saved quantities. The XML files can be accessed in [13].
@@ 3.1 Traditional Correlations
Traditional CHF correlations, such as those by Bowring [14] and Biasi [15], were initially used to predict the collected dataset and were evaluated separately for uniform and non-uniform heating conditions. For uniform cases, Bowring’s correlation yields a root mean square error (RMSE) of 27%, which increases significantly to 74% for non-uniform cases. Biasi’s correlation performs worse, with an RMSE of 62% for uniform cases and 74% for non-uniform ones. Both models occasionally produce unphysical (negative) CHF values, particularly near the limits of their applicability, and exhibit a systematic bias in predicting the CHF location under non-uniform heating. In contrast, the LUT [5], including a non-uniform axial heat flux correction factor [1], demonstrates better overall performance, achieving a lower RMSE of approximately 20% for uniform cases and maintaining greater robustness under non-uniform conditions, with an RMSE of 36%, see Fig. 2.
@@ 3.2 Neural Network Modeling
Number of layers and neurons 5 – 61 – 51 – 28 – 39 – 26 – 21 – 20 – 14 – 1
@@ 5 Acknowledgments
This work was conducted within ME 575 Interdisciplinary Industrial Design Project II at Kansas State University, supported by Westinghouse Electric Sweden. The authors would like to thank Dr. Mooneon Lee and Dr. Cheol Park of the Korea Atomic Energy Research Institute (KAERI) for providing access to the report KAERI/TR-1665/2000.

### v2

# Aggregation of Published Non-Uniform Axial Power Data for Phase II of the OECD/NEA AI/ML Critical Heat Flux Benchmark

@@
Aggregation of Published Non-Uniform Axial Power Data for Phase II of the OECD/NEA AI/ML
Critical Heat Flux Benchmark
@@ 1 Introduction
Critical heat flux (CHF) is a key thermal‐hydraulic phenomenon in which the two‐phase coolant’s heat‐transfer capability drops abruptly, establishing the thermal‐design limit, defined as the departure from nucleate boiling (DNB) in pressurized water reactors (PWRs) and as the critical‐power condition in boiling water reactors (BWRs) [1]. Therefore, accurate prediction of CHF is essential for setting safe operating margins and maximizing reactor power output without risking fuel damage. However, CHF modeling remains challenging due to the complex coupling of local two‐phase hydrodynamics, surface characteristics, transient thermal boundary‐layer development, and geometric influences, factors that vary widely with pressure, flow rate, and heating profile [2, 3].
Traditional CHF correlations often rely on simplified mechanistic assumptions, empirical parametric relationships or look‐up tables (LUT), which can limit their accuracy and adaptability under novel or strongly non-uniform heating conditions [4, 5]. Machine learning (ML) methods, by contrast, can learn high‐dimensional patterns directly from large datasets and have demonstrated improved accuracy and robustness in CHF prediction [6]. However, many ML‐based CHF studies have been carried out independently as proof‐of‐concept demonstrations on limited or proprietary datasets, underscoring the need for coordinated benchmarking efforts. To address this, the Organization for Economic Co-operation and Development / Nuclear Energy Agency (OECD/NEA) Task Force on Artificial Intelligence and Machine Learning for Scientific Computing in Nuclear Engineering organized a CHF benchmark to provide common datasets, blind‐test partitions, and reproducible evaluation protocols for evaluating ML‐based CHF predictors [7].
Phase I of the CHF benchmark, as specified in [7], comprised four tasks: (1) dimensionality analysis for feature selection and extraction; (2) development and optimization of ML‐based CHF regression models based on the Nuclear Regulatory Commision (NRC) database; (3) rigorous model evaluation and uncertainty quantification (UQ); and (4) independent blind testing on withheld data. While Phase I demonstrated the feasibility of AI/ML approaches for CHF prediction under uniform heating, it remained limited by the absence of spatially varying power profiles. Phase II will extend the CHF benchmark to include more complex geometries and boundary conditions, to be used along with transfer learning techniques, advanced UQ, and design optimization, all of which require a richer, more diverse CHF dataset.
To support these Phase II tasks, Kansas State University, in collaboration with Westinghouse Electric Sweden, collected additional CHF measurements in tubes with uniform and non-uniform axial power distributions within the ME 535 “Interdisciplinary Industrial Design Project II” course. This paper reports on those activities and describes the newly digitized uniform and non-uniform CHF datasets; more details can be found in the report [8].
@@ 2.1 Data Source and Characteristics
The primary source of the experimental data collected comes from the Korea Atomic Energy Research Institute (KAERI) technical report [9]. The report investigates how axial heat flux distribution affects the occurrence of CHF in boiling systems. The study includes an extensive set of experimental CHF measurements under both uniform and non-uniform heating conditions, providing valuable data for understanding and modeling CHF behavior in nuclear thermal-hydraulic systems.
Overall, $1539$ CHF test cases were collected/digitized from [9], comprising $651$ uniform heating measurements and $888$ non-uniform heating cases, summarized in Table 1. The experimental setups used water coolant in vertical round tubes with internal diameter of 5.445.4428.328.3 mm and heated lengths of 0.0610.0617.07.0 m.
@@
Equilibrium
quality at CHF
Uniform 0.43 – 18 335.0 – 9561.9 0.582-0.582 – 0.901
Non-uniform 0.43 – 18 328.2 – 8916.0 0.952-0.952 – 0.924
@@ 2.2 Data Collection Workflow
    Energy balance check: The integrated axial flux was compared to published total power; cases with >>2% discrepancy were reprocessed.
@@
For the uniform cases, inlet subcooling and outlet quality are reported in the KAERI report [9]. The remaining inlet and outlet parameters were calculated using IAPWS thermodynamic properties [12]. Unlike the non-uniform cases, the uniform data did not use a detailed 40-node mesh; instead, only inlet and outlet values of local power and temperature/equilibrium quality are recorded. In the non-uniform dataset, the inlet temperature is provided and, combined with the specified heat flux distribution, was used to compute all thermodynamic parameters along the heated tube on a uniform 40-node mesh.
@@ 2.3 Dataset Structure
The quantities were saved in separate (uniform and non-uniform) XML files which are easily readable by computer scripts, see Table 2 that describes the saved quantities. The open-source XML files can be accessed in [13]. Such database can be used to develop and validate predictive CHF models, in particular with respect to the effect of axial power distribution.
@@ 3.1 Traditional Correlations
We first tested traditional CHF correlations, such as those by Bowring [14]:
$q^{\prime\prime}_{\text{cr}}=\dfrac{A-B\,h_{\text{fg}}\,x}{C},$
and Biasi [15]:
$q^{\prime\prime}_{\text{cr}}=15.048\times 10^{7}(100\,D)^{-n}\,G^{-0.6}H\,(1-x),$
where $A$, $B$, $C$ and $H$ are empirical parameters that depend on the thermodynamic state, flow conditions and geometric factors, $h_{\text{fg}}$ is the latent heat of vaporization, $x$ is the steam quality, $D$ is the hydraulic diameter, $G$ is the mass flow rate.
These correlations were used to predict the collected dataset and were evaluated separately for uniform and non-uniform heating conditions. For uniform cases, Bowring’s correlation yields a root mean square error (RMSE) of 27%, which increases significantly to 74% for non-uniform cases. Biasi’s correlation performs worse, with an RMSE of 62% for uniform cases and 74% for non-uniform ones. Both models occasionally produce unphysical (negative) CHF values, particularly near the limits of their applicability, and exhibit a systematic bias in predicting the CHF location under non-uniform heating. In contrast, the LUT [5], including a non-uniform axial heat flux correction factor [1], demonstrates better overall performance, achieving a lower RMSE of approximately 20% for uniform cases and maintaining greater robustness under non-uniform conditions, with an RMSE of 36%, see Fig. 2.
@@ 3.2 Neural Network Modeling
Number of layers
and neurons
5 – 61 – 51 – 28 – 39 –
26 – 21 – 20 – 14 – 1
@@ 5 Acknowledgments
This work was conducted within ME 575 “Interdisciplinary Industrial Design Project II” at Kansas State University, supported by Westinghouse Electric Sweden. The authors would like to thank Dr. Mooneon Lee and Dr. Cheol Park of the KAERI for providing access to the report KAERI/TR-1665/2000.